find (x,y)
y=1/3x + 5/2
x-4y=-8
I keep getting (-6, 9/2)
But the answer sheet shows (9, 4)....

Answer by mananth(16946)

y=1/3x + 5/2
multiply by 6
6y=2x+15
2x-6y=-15...........................(1)
x-4y=-8.............................(2)
solve (1) & (2)

2 x -6 y = -15 .............1
Total value
1 x -4 y = -8 .............2
Eliminate y
multiply (1)by 2
Multiply (2) by -3
4 x -12 y = -30
-3 x + 12 y = 24
Add the two equations
1 x = -6
/ 1
x = -6
plug value of x in (1)
2 x -6 y = -15
-12 -6 y = -15
-6 y = -15 + 12
-6 y = -3
y = 1/2